Decision on Transfer Pricing

Materials used by the Instrument Division of Ziegler Inc. are currently purchased from outside suppliers at a cost of $447 per unit. However, the same materials are available from the Components Division. The Components Division has unused capacity and can produce the materials needed by the Instrument Division at a variable cost of $371 per unit.

a. If a transfer price of $407 per unit is established and 34,800 units of materials are transferred, with no reduction in the Components Division's current sales, how much would Ziegler Inc.’s total operating income increase?
$

b. How much would the Instrument Division’s operating income increase?
$

c. How much would the Components Division's operating income increase?
$
Answer #1
a
Ziegler Inc.’s total operating income increase 2644800 =34800*(447-371)
b
Instrument Division’s operating income increase 1392000 =34800*(447-407)
c
Components Division's operating income increase 1252800 =34800*(407-371)